OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (anonymous):

y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7

OpenStudy (anonymous):

y= 4x+11 Plug in -3 for x and see which one has a answer for -1, they all have a slope of 4

OpenStudy (anonymous):

but B is your answer, a medal would be appreciated and a fan